How many liters of water must be added to 20 L of a 36% soap solution to make it a 12% soap solutio?

Answer by richwmiller(17219)

We have 20 liter of 36% soap
We want to add some 0% soap (water) to make 12% soap
0*x+0.36*20=0.12(20+x)
0*x+7.2=2.4+0.12x
0x-0.12x=2.4-7.2
-0.12x=-4.8
x=40 liter at 0% (0) is added to 20 liter at 36%(7.2) to make 12% of 60 liter(7.2)
check
0*40+0.36*20=0.12(20+40)
0+7.2=0.12(60)
7.2=7.2
ok